MOM Report Paul Soler MICE Operations Manager University of Glasgow MICO April 27, 2009.
Published byModified over 5 years ago
Presentation on theme: "MOM Report Paul Soler MICE Operations Manager University of Glasgow MICO April 27, 2009."— Presentation transcript:
MOM Report Paul Soler MICE Operations Manager University of Glasgow MICO April 27, 2009
Commissioning Needs-Soler2 Commissioning Needs Working Decay Solenoid – Powered up to 900 A, 24/04/09 Many thanks to Matt Hills, Jim Rochford, Mike Courthold, Craig MacWaters, James Leaver, Willie Spensley, Linda… —Cooldown completed Thursday 16 April, temperature ~4.5K —Fri 17-20 April, quench detector (num 5) still not working. —One vacuum pump remains off (tripped 18 April) —Mon 20 April, ramped up to 400 A and Decay Solenoid coils go superconducting (first indication Multi-Layer Insulation working) —Between 21-23 April: o Quench Detector (QD) problem diagnosed to broken wire within QD chassis o 2A fuse associated with QD external relay chain blew and replaced with 5A fuse (need to restore original fuse tolerance) —Fri 24 April: Decay Solenoid powered up in stages to 900 A, then brought down to 870 A (5T B-field), where it remained for 1 hour before powering down once more. —Further tests to be carried out this week. Meeting 28 April to decide procedure for signing off decay solenoid for standard operations
April 27, 2009Commissioning Needs-Soler3 DS Ramp Up 24/04/09 870 A* * Current badly calibrated in control system: 721A in plot corresponds to 870 A real current Total voltage all coils dI/dt ~ 1 hour Peak=900 A
April 27, 2009Commissioning Needs-Soler4 Commissioning Needs Working target – Stands.. —one in R78 for demo —one in ISIS for data Detectors – —Tracker cryostats (Alan Bross at RAL last week) o All pressure release valves tested o Installed emergency shutoff valves on cryostats – all 4 installed and working o Installed exhaust filter o CRYO 3 & 4 now working stably and at low temp (~6K) —CKOV o Need network connection and power in the DSA – info from new temp and humidity sensors not yet logged o Spare HV power supply that died still not replaced Air conditioning —Ducting for A/C venting well underway towards 6 holes in West Wall —A/C unit placed on South Wall mezzanine MICE-EB decided to delay removing synchrotron shielding until August shut-down. Upstream Decay Solenoid shielding reinstated
April 27, 2009Commissioning Needs-Soler5 Issues Documentation inventory/overhaul - ongoing —Instructions for shifters – old documentation needs updating —Safety – need all relevant documents in one obvious place Computing —Need network connections to progress in Hall —Still cannot print on OKI printer in R76 —Wireless in R76 still not installed but network people aware of problem —Need Windows PC in MLCR – still pending —Need connection to GRID from MLCR – pending Need beamline magnets on to test alarm handler —Water will be restored to magnets –this week?